OpenStudy (anonymous):

Find an equation in slope-intercept form (where possible) for the line. Through (0, -5), m = 9/2

OpenStudy (anonymous):

Use your linear function equation: f(x) = mx + b m = 9/2, so f(x) = 9/2(x) + b Sub in (0,-5) and solve for b to find your equation.

OpenStudy (anonymous):

y=9/2x-5 is that right?:

OpenStudy (anonymous):

use point-slope form... y - y0 = m*(x - x0) where (x0, y0) is the given point...